# GIM: A Million-scale Benchmark for Generative Image Manipulation Detection and Localization

### v1

@@ 4.2 Frequency-Spatial Block
When degradation operations are applied, artifacts in manipulated images are tricky to perceive. To improve the local expressive ability and effectively harvest discriminative cues in manipulated images, we design a Frequency-Spatial Bloc(FSB) to extract forgery features in the frequency and spatial domains.
@@
Inspired by the recent work [58, 49, 37, 67], as shown in Figure 3, FSB consists of two branches: a frequency branch and a spatial branch. In the frequency branch, the input $X$ is converted into the frequency domain $\mathcal{F}_{T}(X)$ using the 2D FFT. A learnable filter $G_{i}$ is multiplied to modulate the spectrum and capture the important frequency information. Subsequently, the inverse FFT is applied to convert the feature back to the spatial domain, resulting in the extraction of frequency-aware features $X_{\text{f}}$. In the spatial branch, the input $X$ is processed through convolutional layers and activated using the LeakyReLU function in a separate network to enhance the expressiveness of the features and obtain refined spatial features $X_{\text{s}}$. Then $X_{\text{f}}$ and $X_{\text{s}}$ are concatenated and passed through convolutional layers and the LeakyReLU activation function to obtain enhanced information, which is then combined with the original input $X$ through element-wise summation.
The total process can be formulated by:
@@
where $\odot$ denotes the Hadamard product , $\mathrm{Conv_{L}}$ denotes convolution with LeakyReLU and $\begin{bmatrix}\cdot\end{bmatrix}$ denotes a concatenation operation.
Figure 4: Subtle traces are left by generative manipulation devoid of apparent frequency or structural inconsistency. It fatally modifies the content and this pixel-level perturbation corrupts the image. ShadowTracer captures the intrinsic pattern of the generative tampering and reconstructs the underlying tampering perturbations as priori traces.
4.3 Multi Windowed Anomalous Modelling Module
Image manipulation causes discrepancies at the pixel level. Genuine pixels are expected to exhibit consistency with neighboring pixels, while manipulated pixels may deviate and display anomalies. Former works [62, 35, 28], explore modeling such local inconsistencies. To effectively capture the pixel-level inconsistency between the manipulated and real region, we introduce the Multi Windowed Anomalous Modelling module (MWAM) to model these differences at multiple scales for fine-grained features.
As shown in Figure 3, given the input feature ${F\in H\times W\times C}$, we calculate the difference between each pixel and surrounding pixels within a local window in two branches by Eq.3.
@@ 4.3 Multi Windowed Anomalous Modelling Module
where $u\in\{a,m\}$ denotes average or maximum branches, $\sigma(F)$ is the standard deviation of F, and $w_{\sigma}$ is a learnable non-negative weight vector of the same length as $\sigma$, $F_{a}^{k}$ and $F_{m}^{k}$ are calculated from the average and maximum values of the $k\times k$ windows in each pixel.
Different sizes $k$ are selected to model the inconsistency at different scales. Then, the obtained $N=3$ different-scale $D^{k}_{a}$ and $D^{k}_{m}$ are concatenated and fed into a convolutional network to obtain an anomaly map $M_{a}$ and $M_{m}$ of the same size as the original input.
Additionally, the anomaly score mask $\hat{\bm{S_{u}}}\in H\times W$ of the feature is computed using Eq.4.
$\hat{\bm{f}_{u}}$ $=\mathrm{DW-Conv}\left(\bm{f}\right),$ (4)
$\hat{\bm{S}_{u}}$ $=\operatorname{Sigmoid}\left(\mathrm{Conv}(C,1)\left(\hat{\bm{f}_{u}}\right)\right),$
where the $\mathrm{DW}{-\operatorname{Conv}}$ means a $3\times 3$ Depth-Wise convolution layer.
The element-wise multiplication between anomaly score $\hat{\bm{S}_{u}}$ and the anomaly map $M_{u}$ capture the anomaly information. Next, we calculate the element-wise summation between the resulting anomaly-aware map and the input feature map $X$ to obtain an anomaly-sensitive feature map. The whole process can be described as:
$\hat{X}=X+\hat{\bm{S}_{a}}\times M_{a}+\hat{\bm{S}_{m}}\times M_{m}$ (5)
Table 4: Benchmarking IMDL models for manipulation detection and localization for varying mix-generator comprehensive capability. The models are trained on the whole GIM training set and test on the correspondence test dataset respectively. The detection Cls.Acc(%) and localization AUC(%) and F1(%) are reported. Params and GFLOPS denote the models’ parameters and computational complexity with Params measured in millions (M). $\ddagger$ indicates that the original paper did not provide the code, we reproduced the corresponding code and evaluated it under the same settings
@@
ManTranet[62] 4.0 1009.7 61.08 37.48 80.83 70.99 49.11 83.29 53.99 33.12 74.94
MVSS-Net[12] 146.9 160.0 56.12 23.17 72.03 61.29 33.12 74.94 49.15 14.11 70.09
SPAN[25] 15.4 30.9 53.15 35.62 79.28 60.01 39.46 81.21 59.15 32.1 73.55
PSCC-Net[40] 4.1 107.3 52.28 31.48 83.85 66.52 53.68 86.37 56.33 41.77 85.8
@@
SegFormer[63] 27.5 41.3 64.25 46.17 83.26 78.11 56.77 88.73 69.29 40.19 84.65
GIMFormer (Ours) 95.9 96.2 70.92 58.61 88.25 83.89 77.31 95.42 76.72 56.25 88.31
4.4 Loss Function
For manipulation detection, we adopt a light-weight backbone in  [57] on the fourth stage feature to calculate the final binary prediction $\hat{y}$. For manipulation localization, we utilize the MLP decoder  [63] as the segmentation head to obtain a predicted mask $\hat{M}$. Given the ground-truth label y and mask M, we train GIMFormer with the following objective function:
${\mathcal{L}}={\mathcal{L}}_{cls}(y,{\hat{y}})+\lambda{\mathcal{L}}_{seg}(M,{\hat{M}}),$ (6)
where both ${\mathcal{L}}_{cls}$ and ${\mathcal{L}}_{seg}$ are binary cross-entropy loss, and $\lambda=1$ is a balancing hyperparameter.
4.5 Implementation Details.
Our approach includes two separate training steps. First, we train the ShadowTracer using a synthesized dataset of ImageNet. This training process follows a similar data generation method as described in Section 4.1. Then, we train the encoder and decoder of the model according to the two settings in GIM, as described in Section 3.4. We train our models on eight V100 GPUs with an initial learning rate (LR) of $6e^{-5}$ which is scheduled by the poly strategy with power 0.9 over 40 epochs. The optimizer is AdamW  [41] with epsilon $1e^{-8}$ weight decay $1e^{-2}$, and the batch size is 4 on each GPU.
Table 5: Benchmarking IMDL models for manipulation detection and localization for varying cross-generator generalization capability. The models are trained on the GIM-SD training set and tested on all the test sets to explore the generalization ability. The detection Cls.Acc(%) and localization AUC(%) and F1(%) are reported. $\ddagger$ indicates that the original paper did not provide the code, we reproduced the corresponding code and evaluated it under the same settings
Method GIM-SD GIM-SD(VOC) GIM-GLIDE GIM-DDNM
@@ 4.4 Loss Function
ManTranet[62] 73.12 43.18 80.18 63.18 27.37 72.74 58.21 24.53 74.61 39.52 16.8 58.31
MVSS-Net[12] 56.12 25.13 82.15 56.30 21.2 73.17 53.12 23.17 73.11 48.12 10.11 50.15
SPAN[25] 52.15 47.32 56.93 52.17 29.86 55.17 50.03 30.14 56.08 44.17 14.30 56.67
PSCC-Net[40] 58.24 48.36 87.29 54.79 39.36 83.80 51.11 29.60 79.76 40.52 4.40 48.52
@@
SegFormer[63] 71.93 53.08 87.11 60.03 28.17 80.97 54.11 21.03 75.12 50.21 4.70 51.33
GIMFormer (Ours) 78.96 61.75 90.61 67.19 50.01 84.01 67.01 39.27 81.02 50.23 19.37 60.12
4.6 Comparison with state-of-the-art methods
@@
We compare our methods with various state-of-the-art IMDL methods  [62, 12, 25, 40, 19, 58] on generative manipulation detection and location. In addition, the vanilla SegFormer (MiT-B2) [63] is also compared, since our method is based on that architecture. As shown in Table 4 and Table 5, these methods are tested on various settings to verify the performance and generalization. Note that some methods are not explicitly designed for image-level detection, in which case we use the maximum of the prediction map as the detection statistic. All methods are immersed in the same implementation details.
Cross-generator generalization capability comparison. Table 5 reports the generalization performance of all the methods mentioned. The results show that GIMFormer outperforms all other methods by a significant margin in both in-domain detection ability and cross-domain generalization ability. For in-domain detection ability, our method efficiently catches the subtle artifacts inherent in generative and accurately localizes them. Meanwhile, other methods may encounter confusion as they attempt to learn specific content, potentially leading to challenges in accurately detecting and localizing generative tampering patterns. For cross-domain generalization ability, GIMFormer achieves state-of-the-art performance. It demonstrates cross-detection capability in detecting manipulation using different generative models, as shown in the results of the GIM-GLIDE and GIM-DDNM testsets. Besides GIMFormer works well on Data generated by the same generator from different distributions, while existing methods are confused and have an obvious performance drop, as shown in the results of the GIM-SD(VOC) testset. The qualitative results for visual comparisons are illustrated in Figure 6.
@@
Mix-generator comprehensive capability comparison Table 5 reports the comprehensive performance of all the methods mentioned. GIMFormer also outperforms all other methods on all the test sets, demonstrating its superior ability to identify generative tampering and its comprehensive performance. Existing methods have a low pixel-level F1 score on this benchmark, which means that tampering areas cannot be accurately identified. The qualitative results for visual comparisons are illustrated in Figure 5.
Figure 5: Qualitative results on GIM of comparing GIMFormer with state-of-the-art methods comprehensive capability. From top to bottom, we show examples from the authentic image, GIM-SD, GIM-GLIDE, GIM-DDNM
@@
Figure 6: Qualitative results on GIM of comparing GIMFormer with state-of-the-art methods generalization capability. From top to bottom, we show examples from GIM-SD, GIM-GLIDE, GIM-DDNM
4.7 Ablation Analysis
@@
Effectiveness of proposed module. We consider a simple baseline proposed in  [63] and gradually integrate new key components. Experiments are carried out on the GIM-GLIDE testset. The quantitative results are listed in Table 7. The result shows that ShadowTracer brings significant improvements to the vanilla baseline. With the MWAM, there is an increase of 6.29% in F1 and 2% in AUC, which indicates that the differential information at multiple scales is crucial for accurate tampering localization. The use of FSB to dynamically harvest complementary frequency and spatial cues improves performance, particularly in detection. The results verify that ShadowTracer, FSB and MWAM effectively improve the performance of the baseline model.
Table 6: Ablation results on GIM-GLIDE test dataset using different variants of GIMFormer. All detection Cls.Acc(%), localization AUC(%), and F1(%) scores are reported.
@@
Baseline 78.11 56.77 88.73
+ST 79.96 67.12 91.99
+ST+WMAM 80.19 73.41 94.01
+ST+WMAM+FSB 83.89 77.31 95.42
Table 7: Generalization Experiments of ShadowTracer in percentage (%). ST denotes ShadowTracer and ST(SD) denotes ShadowTracer trained on data generated by Stable-Diffusion.
@@
GIMFormer w ST(SD) 83.1 78.1 95.4 77.4 58.8 89.4
GIMFormer w/o ST 80.0 68.9 91.3 73.1 51.3 86.1
@@
Generalization of ShadowTracer. We initiate the training process of ShadowTracer using data generated by Stable Diffusion and subsequently hold its weights fixed. Following this pretraining phase, we proceed to train the backbone on both the GIM-GLIDE and GIM-DDNM trainsets, with and without the incorporation of ShadowTracer. The detection result Cls.Acc and location results in F1 and AUC are presented in Table 7, revealing that leveraging the pretrained ShadowTracer significantly enhances performance in cross-generator IMDL tasks. Notably, in the GIM-GLIDE dataset with the same manipulation type, leveraging ShadowTracer leads to a remarkable 9% enhancement in F1 and a 4% increase in AUC for location results. Additionally, the benefits extend to the detection aspect, manifesting as a 3% increase in accuracy. In the GIM-DDNM dataset with various manipulation types, there is a 7% enhancement in F1, a 3% increase in AUC for location results and a 4 % increase in accuracy.
5 Conclusion
@@
We address the challenge of detecting and locating generative-based manipulation and provide a reliable database (GIM) for AIGC security. The proposed dataset leverages multiple mainstream generators and tampering methods to provide a variety of generative manipulation data. Additionally, We introduce GIMFormer, a novel transformer-based IMDL framework. ShadowTracer is designed to catch subtle artifacts in generative tampering. While the Frequency-Spatial Block gathers frequency and spatial information, the Multi Windowed Anomalous Modelling module captures pixel-level inconsistencies at multiple scales for fine-grained features. Extensive experiments demonstrate the superior performance of our model, which achieves SoTA results.
Limations. The GIM dataset shares classes with ImageNet and VOC, but it may not encompass future emerging objects due to the evolving variety in the real world. Fine-tuning pre-trained models on new object data can address this issue. Current research primarily centers on image manipulation. The emergence of realistic video generation models like SORA [5] presents fresh challenges as AI manipulation extends into videos. Future plans involve expanding research to address the complexities of video manipulation.
Supplementary Document
@@
In this supplementary document, we include many details of our work: (1) The details of the proposed GIM Datset and compare with related work (Sec. 6) (2) The details of Architecture and implementation (Sec. 7). (3) The additional ablation analysis of GIMFormer(Sec. 8) (4) More Visualizations on GIM dataset (Sec. 9). (5)The societal impact of our work (Sec. 10). (6) The ethics statement of our work (Sec.11).
6 GIM Dataset
6.1 GIM Dataset Configuration

### v2

@@ Introduction
When degradation operations are applied, artifacts in manipulated images are tricky to perceive. To improve the local expressive and harvest discriminative cues in manipulated images, we design a Frequency-Spatial Block (FSB) to extract forgery features in the frequency and spatial domains.
@@
Inspired by the recent work (Rao et al. 2021; Lee-Thorp et al. 2021; Zhang et al. 2022), FSB consists of two branches: a frequency branch and a spatial branch as shown in Figure 3. In the frequency branch, the input $X$ is converted into the frequency domain $\mathcal{F}_{T}(X)$ using the 2D FFT. A learnable filter $G_{i}$ is multiplied to modulate the spectrum and capture the frequency information. Subsequently, the inverse FFT is applied to convert the feature back to the spatial domain, resulting in the extraction of frequency-aware features $X_{\text{f}}$. In the spatial branch, the input $X$ is processed through convolution layers and LeakyReLU function to enhance the expressiveness of the features and obtain refined spatial features $X_{\text{s}}$. Then $X_{\text{f}}$ and $X_{\text{s}}$ are concatenated and passed through convolution layers and the LeakyReLU function to obtain enhanced information, which is then combined with the original input $X$ through element-wise summation. The total process can be formulated by:
@@
where $\odot$ denotes the Hadamard product , $\mathrm{Conv_{L}}$ denotes convolution with LeakyReLU and $\begin{bmatrix}\cdot\end{bmatrix}$ denotes concatenation.
Figure 4: Generative manipulation leaves subtle traces, ShadowTracer identifies intrinsic patterns and reconstructs underlying tampering perturbations.
Multi Windowed Anomalous Modelling Module
Image manipulation causes discrepancies at the pixel level. Genuine pixels are expected to exhibit consistency with neighboring pixels, while manipulated pixels may deviate and display anomalies. Motivated by previous works (Wu, AbdAlmageed, and Natarajan 2019; Kong et al. 2023) that explore local inconsistencies. To effectively capture the pixel-level inconsistency between the manipulated and real region, we introduce the Multi Windowed Anomalous Modelling (MWAM) module to model these differences at multiple scales for fine-grained features.
As shown in Figure 3, for input feature ${F\in H\times W\times C}$, we calculate the difference between pixel and its surroundings within a local window in two branches via Eq.3.
@@
where $u\in\{a,m\}$ denotes average or maximum branches, $\sigma(F)$ is the standard deviation of $F$, and $w_{\sigma}$ is a learnable non-negative weight vector of the same length as $\sigma$, $F_{a}^{k}$ and $F_{m}^{k}$ are calculated from the average and maximum values of the $k\times k$ windows in each pixel. Different sizes $k$ are selected to model the inconsistency at different scales. Then, the obtained $N=3$ different-scale $D^{k}_{a}$ and $D^{k}_{m}$ are concatenated and fed into a convolutional network to obtain an anomaly map $M_{a}$ and $M_{m}$ of the same size as the original input. Additionally, the anomaly score mask $\hat{\boldsymbol{S_{u}}}\in H\times W$ of the feature is computed using.
$\hat{\boldsymbol{f}_{u}}$ $=\mathrm{DConv}\left(\boldsymbol{f}\right),$ (4)
$\hat{\boldsymbol{S}_{u}}$ $=\operatorname{Sigmoid}\left(\mathrm{Conv}(C,1)\left(\hat{\boldsymbol{f}_{u}}\right)\right),$
where the $\mathrm{D}{\operatorname{Conv}}$ means a $3\times 3$ Depth-Wise convolution layer. The element-wise multiplication between anomaly score $\hat{\boldsymbol{S}_{u}}$ and the anomaly map $M_{u}$ capture the anomaly information. Next, we calculate the element-wise summation between the resulting anomaly-aware map and the input feature map $X$ to obtain an anomaly-sensitive feature map. The whole process can be described as:
$\hat{X}=X+\hat{\boldsymbol{S}_{a}}\times M_{a}+\hat{\boldsymbol{S}_{m}}\times M_{m}$ (5)
@@
ManTranet(Wu, AbdAlmageed, and Natarajan 2019) 4.0 1009.7 61.1 37.5 80.8 71.0 49.1 83.3 54.0 33.1 74.9
MVSS-Net(Dong et al. 2022) 146.9 160.0 56.1 23.2 72.0 61.3 33.1 74.9 49.2 14.1 70.1
SPAN(Hu et al. 2020) 15.4 30.9 53.2 35.6 79.3 60.0 39.5 81.2 59.2 32.1 73.6
PSCC-Net(Liu et al. 2022) 4.1 107.3 52.3 31.5 83.8 66.5 53.7 86.4 56.3 41.8 85.8
@@
IML-VIT(Ma et al. 2024) 91.0 136.0 65.2 53.9 84.3 71.3 69.3 92.1 71.2 53.9 85.7
SegFormer(Xie et al. 2021) 27.5 41.3 64.3 46.2 83.3 78.1 56.8 88.7 69.3 40.2 84.6
GIMFormer (Ours) 95.9 96.2 70.9 58.6 88.2 83.9 77.3 95.42 76.7 56.3 88.3
Table 4: Benchmarking IMDL models to evaluate performance. Cls.Acc(%), AUC(%) and F1(%) Params(M) are reported.
Method GIM-SD GIM-VOC GIM-GLIDE GIM-DDNM
@@
ManTranet(Wu, AbdAlmageed, and Natarajan 2019) 73.1 43.2 80.2 63.2 27.4 72.7 58.2 24.5 74.6 39.5 16.8 58.3
MVSS-Net(Dong et al. 2022) 56.1 25.1 82.2 56.3 21.2 73.2 53.2 23.17 73.1 48.1 10.1 50.2
SPAN(Hu et al. 2020) 52.2 47.3 56.9 52.2 29.9 55.2 50.0 30.1 56.1 44.2 14.3 56.7
PSCC-Net(Liu et al. 2022) 58.2 48.4 87.3 54.8 39.4 83.8 51.1 29.6 79.8 40.5 4.4 48.5
@@
IML-VIT(Ma et al. 2024) 77.1 58.1 88.9 54.3 47.2 83.2 52.1 23.1 71.1 50.3 7.2 52.3
SegFormer(Xie et al. 2021) 71.9 53.1 87.1 60.0 28.2 81.0 54.1 21.0 75.1 50.2 4.7 51.3
GIMFormer (Ours) 78.9 61.7 90.6 67.2 50.0 84.0 67.0 39.3 81.0 52.3 19.4 60.1
Table 5: Benchmarking IMDL models to evaluate generalization. Cls.Acc(%), AUC(%) and F1(%) are reported.
Loss Function
For detection, we adopt a lightweight backbone in  (Wang et al. 2020) on the fourth stage feature for binary prediction $\hat{y}$. For localization, we utilize the MLP decoder  (Xie et al. 2021) as the segmentation head to obtain a predicted mask $\hat{M}$. Given the ground-truth label $y$ and mask $M$, we train GIMFormer with the following objective function:
${\mathcal{L}}={\mathcal{L}}_{cls}(y,{\hat{y}})+{\mathcal{L}}_{seg}(M,{\hat{M}}),$ (6)
where both ${\mathcal{L}}_{cls}$ and ${\mathcal{L}}_{seg}$ are binary cross-entropy loss.
Implementation Details
Our approach includes two separate training steps. First, we train the ShadowTracer using the dataset generated from ImageNet. This training process follows a similar data generation method as mentioned in the former chapter. Then, we train the encoder and decoder of the model according to the two settings in GIM, as described in the previous section. We train our models on 8 V100 GPUs with an initial learning rate of $6e^{-5}$ which is scheduled by the poly strategy with power 0.9 over 20 epochs. The optimizer is AdamW  (Loshchilov and Hutter 2017) with epsilon $1e^{-8}$ weight decay $1e^{-2}$, and the batch size is 4 on each GPU.
Comparison with state-of-the-art methods
@@
We compare our methods with various state-of-the-art IMDL methods 111 $\ddagger$ indicates that the original paper does not provide the code, we reproduce the code and evaluate it under the same settings. and the vanilla SegFormer (MiT-B2) (Xie et al. 2021). As shown in Table 4 and Table 5, these methods are tested in two settings to evaluate the performance. Note that some methods are not explicitly designed for image-level detection, in which case we use the maximum of the prediction map as the detection statistic. All methods are immersed in the same implementation details.  More experiments are included in the Appendix.
@@
Mix-generator comparison. Table 4 reports the performance of all methods. GIMFormer outperforms other methods, demonstrating its superior ability to identify generative manipulation. Existing methods exhibit low pixel-level F1 scores on this benchmark, indicating that tampering areas are not accurately identified. The qualitative results for visual comparisons are illustrated in Figure 5.
Cross-generator comparison. Table 5 reports the generalization of all the methods. The results show that GIMFormer outperforms all other methods in both in-domain and cross-domain. For in-domain experiments, our method catches the subtle artifacts inherent in manipulation and accurately localizes them. Other methods may encounter confusion as they attempt to learn specific content, potentially leading to challenges in accurately detecting and localizing generative tampering patterns. For cross-domain experiments, GIMFormer demonstrates well generalization in detecting manipulation using different generative models, as shown in the results of the GIM-GLIDE and GIM-DDNM testsets. Besides GIMFormer works well on data generated by the same generator from different distributions, while existing methods have an obvious performance drop, as shown in the results of the GIM-VOC testset. The qualitative results for visual comparisons are illustrated in Figure 6.
@@
Figure 5: Qualitative results on GIM of comparing GIMFormer with state-of-the-art methods performance.
Figure 6: Qualitative results on GIM of comparing GIMFormer with state-of-the-art methods generalization capability.
Ablation Analysis
@@
Effectiveness of the proposed module. We consider a simple baseline proposed in  (Xie et al. 2021) and gradually integrate new key components. Experiments are carried out on the GIM-GLIDE testset. The quantitative results are listed in Table 6. The result shows that ShadowTracer brings significant improvements to the vanilla baseline. With the MWAM, there is an increase of 6.29% in F1 and 2% in AUC, which indicates that the differential information at multiple scales is crucial for accurate tampering localization. The use of FSB to dynamically harvest complementary frequency and spatial cues improves performance, particularly in detection. The results verify that ShadowTracer, FSB and MWAM effectively improve the performance of the baseline model.
@@
Baseline 78.1 56.8 88.7
+ShadowTracer 79.9 67.1 92.0
+ShadowTracer+WMAM 80.2 73.4 94.0
+ShadowTracer+WMAM+FSB 83.9 77.3 95.4
Table 6: Ablation results of GIMFormer variants. Cls.Acc(%), AUC(%) and F1(%) scores are reported.
@@
GIMFormer w ShadowTracer 83.1 78.1 95.4 77.4 58.8 89.4
GIMFormer w/o ShadowTracer 80.0 68.9 91.3 73.1 51.3 86.1
Table 7: Generalization Experiments of ShadowTracer. Cls.Acc(%), AUC(%) and F1(%) scores are reported.
@@
Generalization of ShadowTracer. To evaluate the generalization of ShadowTracer, we train ShadowTracer using data generated by Stable Diffusion and subsequently hold its weights fixed. We proceed to train the backbone on the GIM-GLIDE and GIM-DDNM trainsets, with and without the incorporation of ShadowTracer. The result in Table 7 reveals that leveraging the pretrained ShadowTracer increases performance in cross-generator IMDL tasks. Across GIM-GLIDE and GIM-DDNM, ShadowTracer achieves up to 9% F1, 4% AUC and 4% accuracy improvements.
Conclusion
@@
We address the challenge of detecting and locating generative manipulation and provide a reliable database GIM for AIGC security. This dataset leverages multiple generators to provide diverse generative manipulation data. Based on this, we design a benchmark for IMDL methods with two settings. We also introduce GIMFormer, a novel transformer-based IMDL framework. Extensive experiments demonstrate that GIMFormer achieves SOTA performance.
Appendix
@@
In this Appendix, we include more details of our work: (1) The details of the proposed GIM Datset and compare with related work (2) Details of the GIMFormer architecture and implementation (3) The additional ablation analysis of GIMFormer (4) More Visualizations on GIM dataset. (5)The societal impact of our work . (6) The ethics statement of our work.
Appendix A GIM Dataset
GIM Dataset Configuration